What is the fifth term of a sequence whose first term is 5 and whose common ratio is 3?

Answer by Euler(11)

Hey there,
okay, now here you have a geometric progression. You know it's one whenever the question states the first term and the common ration. So what you do now is, plug in the values and get your answer! :)
Formula:
a(n)=a*r^(n-1)
where,
a=first term
r=common ratio
n=the term
a(5)=5*3(5-1)
a(5)=5*81
a(5)=405